OpenStudy (anonymous):

the perimeter of a rectangle park is 400 feet. the length of the park is 40 feet longer than the width. find the length of the park

OpenStudy (blurbendy):

2x+2(x+40)=400 2x + 2x + 80 = 400 4x + 80 = 400 4x = 320 x = 80 ft